P1. 5pts Use De Morgans law to compute complements of the following Boolean expressions: (a) A: C + Ā:B:C (b) Ā: B.C.+ A:B .Can I get this problems solved please, I need them really worked correct. Thanks
Answer #1

@ ACT Ā B C = (AC) (ABC) = (a + c) (A+B+C) = Ã. A+ ĀBTĀCTAC+ BCT c.c = ĀB + C (ATĀJ + BC tł = ĀB + C + B cao = ĀB + C (+1+B)= (B+ActÃE TĀD+CD) (A + B5 + B C + CD +BD) (AB + ACT AD + BC + BD + BC + čo + BD + CD) A. @ ABC TAŘ = (ABC) (AR) = = (ABC) (ĀTĀCD TAD = AB + BCD BCD TAC +Ā DĀ B CTABD + & AC + D + ĀCD +ĀCDTÃD. - AB + BCD + ĀC (1+B+D) + ĀDT ĀRD T ĀCD+ÃO = AB + BCD + Ãa (A + B + C). (A+B+C). (A+C) Multiply (A + B +C) [A + AC + AB + BC + AC + C, 2] ciero - At AC + ABT ABC & AT +AR+ABC + ABC+- ABC TĀ BD + Co - AB CE+ og tao ☺ (A.D+B), (B+C+D) ABD & ACD+AD+ B C + BD ADCB + C +1] + BC + RD AD+BE+BD. = 1 AD + B [E+D